sir please tell what is the net resistance between points P and Q.

Grade:10

7 Answers

Archita
8 years ago
1/net resistance =1/R +1/R
                         = 2/R
 
net resistance = R/2
                 proided your pand q is located in the point as shown in your attachement.
manoj singh
8 years ago
the answer is
Archita
8 years ago
what did u mean?
manoj singh
8 years ago
the answer is 2R/5 iwant the solution how to find it. 
Prajwal Kavad
8 years ago
total resistance in 5R/2 and not 2R/5 it is simple series and parallel circuit plz refer your scool textbooks it will have enough theory to sove this problem
Prajwal Kavad
8 years ago
and between p and q its just R/2
Ramakrishna Manduri
8 years ago
all ur ans r wrong and the ans is none